OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax, and x represents the original price of the item. C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale? Numerical Answers Expected!

OpenStudy (anonymous):

C = 0.765x + .06(0.765x) C = 0.765(100) + .06(0.765(100) C = 76.5 + .06(76.5) C = 76.5 + 4.59 C = 81.09 @jim_thompson5910 ......can you take a peek at this......do I still need to take off the discounts......or did I do it correctly by just following the equation ?

jimthompson5910 (jim_thompson5910):

the discounts were already calculated and applied when they came up with this equation

jimthompson5910 (jim_thompson5910):

so that saves you from having to figure out the overall discount percentage or anything like that which means all you have to do is plug in x = 100 and evaluate

OpenStudy (anonymous):

ok.....I understand......thank you

jimthompson5910 (jim_thompson5910):

you're welcome